OpenStudy (anonymous):

find the height in feet of the rock after 10 seconds in the air

OpenStudy (lizz123):

Is there more information?

OpenStudy (anonymous):

yes it the whole problem do you want me to write it out

OpenStudy (lizz123):

sure :p

OpenStudy (anonymous):

that is it

OpenStudy (lizz123):

that website says, "AccessDeniedAccess"

OpenStudy (anonymous):

cant see that its restricted

OpenStudy (anonymous):

what i think you are trying to do is have a rock fall to the ground?

OpenStudy (lizz123):

that website says, "AccessDeniedAccess"

OpenStudy (anonymous):

d = Velocity(time) + (1/2)(9.8)time^2

OpenStudy (anonymous):

a rock is launched from a cannon height ,h(x), can be represented by a quadratic function in terms of x in seconds after 1 second the rock is 58 feet in the air after 2 seconds it is 112 feet in the air

OpenStudy (anonymous):

is this a math class or a physics?

OpenStudy (anonymous):

math

OpenStudy (anonymous):

so we need to make a quadratic equation such as y = x^2 that fits that data

OpenStudy (anonymous):

ok so what i do next

OpenStudy (anonymous):

https://www.youtube.com/watch?v=DHYPjhQNHIo

OpenStudy (anonymous):

we have two points so we can use this formula

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

points = (1,58) (2,112)

OpenStudy (anonymous):

so how do i do this

OpenStudy (anonymous):

sorry i suck at math and i really need to pass this test

OpenStudy (lizz123):

y = vi.t + 0.5.g.t^2 58 = vi + 0.5*32 vi = 42 ft/s y10 = 42*10 + 0.5*32*10^2 = 2000'

OpenStudy (anonymous):

is that the answer

OpenStudy (anonymous):

seems like we need more information. otherwise we would just have to use physics formulas

OpenStudy (anonymous):

ok cool

OpenStudy (lizz123):

that is what i used to get the answer

OpenStudy (anonymous):

ok

OpenStudy (lizz123):

so yeah 2000' is the answer

OpenStudy (anonymous):

it said it was wrong